OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p) p = 6 p = 5 p = 7 p = 12

jhonyy9 (jhonyy9):

-6p+7=6p-9+40-16p 4p=24 p=24/4 =6 p=6